OpenStudy (anonymous):

find formulae for fogoho ? f(x)=x power 2 + 1 , g(x)=1/x , h(x)=x power 3

OpenStudy (anonymous):

\[ (f\circ g\circ h)(x) = f(g(h(x))) = f(g(x^3))=f\left(\frac{1}{x^3}\right) = \left(\frac{1}{x^3}\right)^2+1 \]

OpenStudy (anonymous):

it is a answer ?

OpenStudy (anonymous):

I dunno.

OpenStudy (anonymous):

Does it looks like the answer?

OpenStudy (anonymous):

no think next step is 1/x power 5 + 1

OpenStudy (anonymous):

what you say?

OpenStudy (anonymous):

How do you get power 5?

OpenStudy (anonymous):

\[ (x^3)^2 = (x^3)(x^3) \]

OpenStudy (anonymous):

because in multiplications powers are add so 3+2 = 5

OpenStudy (anonymous):

Yeah, but you're multiply \(x^3\) and \(x^3\). So it is 3+3

OpenStudy (anonymous):

no i think its x power 5

OpenStudy (anonymous):

Okay do you agree with this: \[ x^a\times x^b = x^{a+b} \]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Do you agree with this: \[ (x^a)^2 = x^a \times x^a \]

OpenStudy (anonymous):

no x power 2a is the answer what you say?

OpenStudy (anonymous):

I say they're the same damn thing.

OpenStudy (anonymous):

If a = 3, then 2a = 6. 2a is not 5

OpenStudy (anonymous):

oh ya right so sorry its my mistake

OpenStudy (anonymous):

next step is 1/x power 6 +1 ?

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

1+ x power 6 / xpower 6 ?

OpenStudy (anonymous):

i think this is the answer ?

OpenStudy (anonymous):

r u there ?

OpenStudy (anonymous):

\[\left(\frac{1}{x^3}\right)^2+1=\frac{1}{x^6}+1=\frac{x^6+1}{x^6}\]

OpenStudy (anonymous):

1+x power 6 / x power 6 is right

OpenStudy (anonymous):

if you mean (1 + x^6)/x^6 then yes :)

OpenStudy (anonymous):

you mean i'm correct ?

OpenStudy (anonymous):

if 1 + x power 6 is on top then yes :)

OpenStudy (anonymous):

thanks